OpenStudy (anonymous):

a line passes through the point (0,2) and has a slope of -1/4 what is the equation of the line

OpenStudy (nikato):

Is there any particular form you want it in?

OpenStudy (anonymous):

i just need the equation

OpenStudy (nikato):

Ok. The fastest would be slope-intercept form. y=mx+b Where m is the slope and b is the y-intercept

OpenStudy (nikato):

Because it passes through (0,2) It intercepts the y-axis at 2 So b=2 And you're given the slope =-1/4 So m=-1/4

OpenStudy (nikato):

Plug them in at your equation is y=-1/4 x +2